Rico's collection of quarters and dimes contains $32.85. There are 171 coins in all. How many quarters and how many dimes are in Rico's collection?
Answer by stanbon(57219)

Rico's collection of quarters and dimes contains $32.85. There are 171 coins in all. How many quarters and how many dimes are in Rico's collection?
----------------------------
Equations:
Quantity Equation: q + d = 171
Value Equation...:25q+10d = 3285 cents
-------------------------
Modify to get:
10q + 10d = 1710
25q + 10d = 3285
-----------------------
Subtract 1st from 2nd to solve for "q":
15q = 1575
q = 105 (# of quarters)
--------------
Substitute into q + d = 171 to solve for "d":
105 + d = 171
d = 66 (# of dimes)
---------------------------
Cheers,
Stan H.